The Alpine House, Inc., is a large retailer of snow skis. The company assembled the information shown below for the quarter ended March 31:

1. Prepare a traditional income statement for the quarter ended March 31.

2. Prepare a contribution format income statement for the quarter ended MarchThe Alpine House, Inc., is a large retailer of snow skis. The company assembled the information shown below for the quarter eTotal fixed administrative expense Beginning merchandise inventory Ending merchandise inventory Merchandise purchases 110,0004 The Alpine House, Inc., is a large retailer of snow skis. The company assembled the information shown below for the quarter 31.

3. What was the contribution margin per unit?
Answer #1

Solution 1:

Units = 989000/430 = 2300 units

Alpine House, Inc.
Traditional format Income Statement
Sales $9,89,000
Less: Cost Of Goods sold:
Beginning Merchandise inventory $80,000
Add: Merchandise purchases $2,90,000
Merchandise available for sale $3,70,000
Less: Ending merchandise Inventory $1,20,000
Cost of Goods sold $2,50,000
Gross Profit $7,39,000
Less: Selling and Administrative Expenses:
Variable Selling Expenses $1,05,800
Variable Administrative Expenses $36,800
Fixed Selling Expenses $1,45,000
Fixed Administrative Expenses $1,10,000
Total expenses $3,97,600
Net Operating Income $3,41,400

Solution 2:

Alpine House, Inc.
Contribution format Income Statement
Sales $9,89,000
Variable Expenses:
Coat of goods sold ($80000+$290000 - $120000) $2,50,000
Variable Selling Expenses $1,05,800
Variable Administrative Expenses $36,800
Total Variable Expenses $3,92,600
Contribution Margin $5,96,400
Fixed Expenses:
Fixed Selling Expenses $1,45,000
Fixed Administrative Expenses $1,10,000
Total Fixed Expenses $2,55,000
Net Operating Income $3,41,400

Solution 3:

Contribution margin per unit = 596400/ 2300 = $259.30